Two tonnes of an iron ore containing 94% of Fe2O3 produces iron inpure state . Calculate the weight of iron

Grade:11

1 Answers

Vikas TU
7 years ago
Two tonnes of an iron ore containing 94% of Fe2O3 =
0.94 * 2 = 1.88 tonnes of Fe20O3
Weight of Fe = 56/160 * 1.88 = 0.666 tonnes approx.
where 56 and 160  are the atomic weights of Fe and Fe2O3 respectively.